# Rethinking the Atmospheric Scattering-driven Attention via Channel and Gamma Correction Priors for Low-Light Image Enhancement

### v1

@@ 4.2 Loss Functions
HDR L1 loss, as introduced in [23], is computed in the tone-mapped domain since HDR (High Dynamic Range) images are typically viewed after tone-mapping. To achieve this, they apply the widely used µ-law function to calculate the loss:
$T=\text{sgn}(x)\frac{\log(1+\mu x)}{\log(1+\mu)}$ (14)
@@
$L_{\text{SSIM}}=1-\text{SSIM}(\hat{Y},Y^{GT})$ (16)
4.3 Datasets and Evaluation Metrics
@@
For evaluation, we apply our approach to both paired and unpaired datasets. For paired data, we use the LOLv1 and LOLv2 datasets [5], benchmarks for the Low-Light Image Enhancement task. LOLv1 includes 485 images for training and testing, while LOLv2 consists of two subsets: real-captured and synthetic. The real-captured subset (LOLv2 Real) has 689 images for training and 100 for testing, while the synthetic subset (LOLv2 Synthetic) has 900 training images and 100 testing images. For unpaired data, we utilize five datasets: LIME [11], MEF [24], NPE [25], VV [26], and DICM [27]. Since these datasets lack ground truth references for paired evaluation, we assess performance using the NIQE metric, which is widely used to evaluate the naturalness of images.
4.4 Evaluation Results
@@
As shown in Fig. 1 and Tables 1 and 2, our approach reaches a high standard compared to others, which ranked third on paired datasets and ranked first on five sets of unpaired data, while the model remains lightweight design with low numbers of parameters and FLOPs. The visualized comparisons are shown in Figs. 6 and 7. Also, we figured out that the approaches with CPGA architecture present better quality on unpaired data, which means that our theoretical equations assumptions for improvement are closer and related to nature, making the image more realistic. Furthermore, our method successfully improves the performance via the same architecture by 5% SSIM compared to the CPGA-Net while maintaining the lightweight design.
@@
Table 1: Comparison to SOTA methods on paired datasets [5]. We represent the first and second ranks with bold and underlined, respectively.
@@
Retinex-Net [5] 16.77 0.425 0.474 18.37 0.723 17.14 0.756 0.555 79.610
KinD [7] 17.65 0.771 0.175 14.74 0.641 17.28 0.758 8.017 63.677
EnGAN [6] 17.54 0.664 0.326 18.23 0.617 16.49 0.771 54.409 22.192
@@
Table 2: Image quality comparison on unpaired data [11, 24, 25, 26, 27]. We represent the first and second ranks with bold and underlined, respectively.
@@
Method MEF LIME NPE VV DICM Avg
Low-light Image 4.265 4.438 4.319 3.535 4.255 4.162
NPE [25] 3.524 3.905 3.953 2.524 3.760 3.533
LIME [11] 3.720 4.155 4.268 2.489 3.846 3.696
EnlightenGAN [6] 3.232 3.719 4.113 2.581 3.570 3.443
KinD [7] 3.883 3.343 3.724 2.321 2.989 3.252
Zero-DCE [13] 4.041 3.789 3.504 2.753 3.102 3.438
RUAS [14] 4.140 4.290 4.871 3.509 4.542 4.270
IAT [15] 3.619 4.172 3.289 2.527 3.033 3.328
PairLIE [16] 4.086 4.311 4.089 3.160 3.242 3.778
FLIGHT-Net [17] 3.549 3.705 3.331 2.944 2.898 3.285
CPGA-Net [8] 3.870 3.707 3.548 2.264 2.693 3.216
CPGA-Net+ (Proposed) 3.500 3.066 3.083 1.913 2.827 2.878
Figure 6: Visual comparison on paired datasets [5].
Figure 7: Visual comparison on unpaired datasets [11, 24, 25, 26, 27].
5 Ablation Study
In this section, we analyze the effectiveness of each systematic module and training technique, including the design of the CPGA block, the number of CP blocks, and loss functions.
5.1 Systematic Design of Plug-in Attention
As shown in Table 3, our method effectively fuses the gamma correction from the global branch to the local branch, resulting in improved overall performance and demonstrating the strength of our approach. By grounding the attention mechanism in gamma correction, we ensure that the enhancement process remains aligned with the non-linearities inherent in both the imaging process and human perception.
Table 3: Ablation study of systematic design.
@@ 5.1 Systematic Design of Plug-in Attention
(a) 18.36 0.743 0.297 0.030 4.78
(b) 20.87 0.803 0.205 0.050 6.929
(c) 22.53 0.812 0.205 0.060 9.356
5.2 The Numbers of Channel-Prior Blocks
We explore how varying the number of CP blocks affects the model’s capacity to enhance image quality. The results, summarized in Table 4, show that increasing the number of CP blocks leads to an improvement from 0 to 2 blocks but show no significant changes from 2 to 4 blocks. However, both the number of parameters and computational cost (FLOPs) increase with more CP blocks, introducing greater computational demands. Therefore, the optimal number of CP blocks should balance performance gains with resource efficiency. For the final design of our approach, we selected 2 CP blocks to achieve a lightweight and efficient design.
Table 4: Ablation study of the number of CP blocks. $N=2$ is the default setting of our approach.
@@ 5.2 The Numbers of Channel-Prior Blocks
5.3 Loss Functions
This section examines the impact of various loss function combinations on the model’s performance. We tested L1 loss, Perceptual loss, HDR L1 loss, and SSIM loss, with the results summarized in Table 5. Using the default settings as in CPGA-Net, the combination of L1 and Perceptual losses performs well, yielding a PSNR improvement of +2.82 dB and an SSIM increase of +2.4%. The HDR L1 loss significantly enhances all three metrics, with PSNR increasing by +2.86 dB, SSIM by +4.1%, and LPIPS decreasing by 0.09. While SSIM loss improves its specific metric with an SSIM boost of +3.1%, it is less effective in enhancing PSNR. Ultimately, combining all these losses results in the best overall performance for supervision, which has improved with PSNR 4.04 dB, SSIM +8.3%, and LPIPS -0.12.
Table 5: Ablation study of loss functions.
@@ 5.3 Loss Functions
(a) 18.49 0.729 0.323
(b) 21.31 (+2.82) 0.753 (+2.4%) 0.250 (-0.07)
(c) 21.35 (+2.86) 0.770 (+4.1%) 0.238 (-0.09)
(d) 20.41 (+1.92) 0.760 (+3.1%) 0.243 (-0.08)
(e) 22.53 (+4.04) 0.812 (+8.3%) 0.205 (-0.12)
6 Conclusion
This work looks deeper at CPGA-Net, utilizing it as an attention mechanism grounded in theoretical formulas. We propose a stacked and modularized attention module to focus on image details. Additionally, we integrate gamma correction into the local branch, creating a Plug-in Attention module for each CP block. This enhancement makes our approach lightweight yet state-of-the-art in performance, maintaining strong efficiency and stable operation across devices with limited computational resources. In the future, we aim to integrate our approach into High Dynamic Range (HDR) imaging and exposure fusion, improving detail preservation in both bright and dark areas by leveraging brightness sensitivity through prior knowledge, such as channel and gamma-correction priors. This will enhance the output’s dynamic range and overall fidelity in real-world applications.
@@ Acknowledgment
This work was partly supported by the Ministry of Science and Technology, Taiwan, under Grant NSTC 113-2221-E-033-055.

### v2

@@ III-C Loss Functions
HDR L1 loss, as introduced in [23] is computed in the tone-mapped domain since HDR (High Dynamic Range) images are typically viewed after tone-mapping. To achieve this, they apply the widely used µ-law function to calculate the loss:
$T=\text{sgn}\frac{\log(1+\mu x)}{\log(1+\mu)}$ (14)
@@
$L_{\text{SSIM}}=1-\text{SSIM}(\hat{Y},Y^{\text{GT}})$ (16)
For unsupervised learning of our approach, the only optimal for brightening is the histogram-equalized low-light image, which is a good aim for the contrast but not for the details. Thus, we utilized total variation loss [24, 25] for denoising and improving the smoothness:
$L_{TV}=\frac{1}{hwc}\sum_{i,j,k}\sqrt{\left(\hat{Y}_{i,j+1,k}-\hat{Y}_{i,j,k}\right)^{2}+\left(\hat{Y}_{i+1,j,k}-\hat{Y}_{i,j,k}\right)^{2}}$ (17)
where $h$, $w$, and $c$ represent the height, width, and number of channels, respectively; $i$, $j$, and $k$ represent the indices corresponding to height, width, and channel, respectively.
III-D Revealing the Explainable Mechanisms Behind Degradation
Our method follows a rule-based learning strategy, leveraging the ATSM and the Retinex theory to perform image processing based on physical models. The ATSM simulates the scattering and absorption of light in uneven media, while the Retinex theory, inspired by retinal imaging, models the human visual system. Although these methods provide solid theoretical support from both physical and physiological perspectives, challenges remain in the interpretability of neural networks. While applying the models to deep learning, it is difficult to explain its decision-making mechanisms due to the highly nonlinear nature of deep learning models. Similarly, these are hard to define and are barely explained through experience or a simple understanding of the equations from the generated features in the neural network. To improve the interpretability of our approach, we utilize the traditional image degradation model, which offers a more explicit theoretical framework for the deep learning process. This effectively enhances the transparency of the entire neural network, opening the “black box” of image enhancement.
Initially, the fundamental image degradation model in matrix form [1] can be defined as:
[…]
where G is the image after degradation, H is the unknown degradation kernel, N denotes additive noise, and F is the image before degradation. The restoration process is aimed at estimating F by:
$\hat{\textbf{F}}=\textbf{H}^{-1}(\textbf{G}-\textbf{N})=\textbf{H}^{-1}\textbf{R}$ (19)
[…]
Figure 6: Visualization of extracted features corresponding to Eq. (3) and (11) [8]. (a) Low-light image $L$; (b) a portion of $L$; (c) $\tilde{t}$ (contribution proportion from $L$); (d) An estimated image of $\tilde{A}$; (e) Reconstructed low-light image $R$; (f) Gamma-corrected image $R^{\gamma}$; (g) Enhanced image $\hat{R}$.
We attempt to further verify our hypotheses for the local branch through least-squares optimization [26]. First, we make the following simple assumptions: we assume that $\hat{A}$ and $L$ can be expressed as the noise-free reconstructed image $R^{\text{GT}}$ with additional composite noises that consist of three-channel mixing factors, lighting changes, and inherent thermal noises, which are particularly prominent in low-light environments, as depicted below:
$\tilde{A}=R^{\text{GT}}+N_{A}$ (20)
$L=R^{\text{GT}}+N_{L}$ (21)
where $L$ is the low-light image, $N_{A}$ and $N_{L}$ denote the noise components associated with $\tilde{A}$ and $L$, respectively. While estimating $N_{A}$, we assume $N_{L}$ is fixed once the low-light image $L$ is given. Then, we can define the following residual error $r(\tilde{t},N_{A})$ by combining Eq. (3), Eq. (20), and Eq. (21):
$r(\tilde{t},N_{A})=R-R^{\text{GT}}=(1-\tilde{t})N_{A}+\tilde{t}N_{L}$ (22)
where $R$ is the reconstructed image, $\tilde{t}$ is a positive parameter used to adjust the reconstruction process. We aim to find $N_{A}$ and $\tilde{t}$ such that the residual error is minimized, allowing $R$ to approach $R^{\text{GT}}$.
We further define a cost function $C(\tilde{t},N_{A})$ for the least-square optimization process:
$C(\tilde{t},N_{A})=\frac{1}{2}\left[r(\tilde{t},N_{A})\right]^{2}$ (23)
Next, to find the stationary points in the optimization process, we conduct partial differentiation of $C(\tilde{t},N_{A})$ with respect to $\tilde{t}$ and $N_{A}$:
$\frac{\partial C}{\partial\tilde{t}}=r(\tilde{t},N_{A})\cdot\left({N_{L}-N_{A}}\right)=0$ (24)
$\frac{\partial C}{\partial N_{A}}=r(\tilde{t},N_{A})\cdot\left(-\tilde{t}\right)=0$ (25)
Case I: $r(\tilde{t},N_{A})\neq 0$: We have
$(N_{A}-N_{L})=0\quad\text{and}\quad\tilde{t}=0$ (26)
which leads to the solution $\tilde{t}=1$ and $N_{L}=N_{A}$, resulting in:
$r_{\text{min}}(\tilde{t},N_{A})=N_{A}$ (27)
However, $\tilde{t}=0$ results in $R=L$ from Eq. (3), meaning that no restoration effect is involved at all (identity mapping from the input $L$ to the output $R$).
Case II: $r(\tilde{t},N_{A})=0$: We have
$R-R^{\text{GT}}=0\implies R=R^{\text{GT}}$ (28)
and
$(1-\tilde{t})N_{A}+\tilde{t}N_{L}=0\implies N_{L}=\left(1-\frac{1}{\tilde{t}}\right)N_{A},\quad\tilde{t}\neq 0$ (29)
Therefore, the optimal solution exists when Eq. (29) holds.
The results show that the enhancement follows the degradation formula, with the ATSM simulating the network’s processing of illumination, reflection, and noise, supporting the rationality of neural networks in LLIE. This enhances the model’s transparency and reveals its interpretability mechanism, offering deeper insights into how neural networks function in image enhancement tasks.
From a learning mechanism perspective, deep learning exhibits significant similarities to human cognition, emphasizing that the most prominent features carry essential information rather than relying on intricate details, with $\tilde{A}$ for detail restoration representing more critical information than $\tilde{t}$ for contrast and saturation. By modularizing and extending the understanding through attention mechanisms, we effectively capture the relationships between local and global information, enhancing image quality in low-light conditions and yielding better results across various complex scenarios. This method underscores the potential for integrating traditional theory with deep learning models, offering valuable insights for future technological advancements and opening new research opportunities in image enhancement.
IV Experiment Results
This section compares our approach with several SOTA methods on benchmark datasets, including paired and unpaired datasets.
IV-A Datasets and Evaluation Metrics
@@
For evaluation, we apply our approach to both paired and unpaired datasets. For paired data, we use the LOLv1 and LOLv2 datasets [5], benchmarks for the LLIE task. LOLv1 includes 485 images for training and testing, while LOLv2 consists of two subsets: real-captured and synthetic. The real-captured subset (LOLv2 Real) has 689 images for training and 100 for testing, while the synthetic subset (LOLv2 Synthetic) has 900 training images and 100 testing images. For unpaired data, we utilize five datasets: LIME [11], MEF [27], NPE [28], VV [29], and DICM [30]. Since these datasets lack ground truth references for paired evaluation, we assess performance using the NIQE metric, which is widely used to evaluate the naturalness of images. For object detection applications, we apply our approach to the Exclusively Dark (ExDark) dataset [31], which consists of 7,363 low-light images from very low-light environments to twilight with 12 object classes and provides a suitable benchmark for evaluating object detection performance in such conditions.
IV-B Implementation Details
Our experiments are conducted and evaluated with an NVIDIA GeForce RTX 3090 GPU. For the preprocessing of training, we cropped the image into $256\times 256$ pixels with random translation. We set 600 epochs for training. An initial learning rate of $10^{-3}$is set for training with the Adam optimization scheme, and the learning rate changing cycle is 67 epochs by the Cosine Annealing scheduler. To address computational demand, we implemented simplified block pruning named “Block Design Simplification (BDSF)” in our lightweight version, directly removing blocks in the local branch for inference. Furthermore, we have also applied our approach as unsupervised learning using histogram equalization, demonstrating its flexibility. When using unsupervised learning, we trained on LOLv2 Real for 50 epochs, which consists of more realistic data and does not require any normal exposure images for training.
IV-C Evaluation Results
@@
As shown in Tables I and II, our approach reaches a high standard compared to others, which ranked third on paired datasets and ranked first on five sets of unpaired data, while the model remains lightweight design with low numbers of parameters and FLOPs. The visualized comparisons are shown in Figs. 7 and 8. Also, we figured out that the approaches with CPGA architecture present better quality on unpaired data, which means that our theoretical equations assumptions for improvement are closer and related to nature, making the image more realistic. Furthermore, our method successfully improves the performance via the same architecture by 5% SSIM compared to the CPGA-Net while maintaining the lightweight design.
For the comparison of unsupervised approaches, as shown in Table II we ranked first compared to other unsupervised approaches with better contrast. This demonstrates the robustness of our theoretical-based network architecture when using simple supervision of histogram-equalized images. However, there are more noticeable defects and distortions due to the lack of strong supervision of the details, as shown in Fig. 9. This will be a focus for our future work.
@@
TABLE I: Comparison to SOTA methods on paired datasets [5]. We represent the first and second ranks with bold and underlined, respectively. BDSF means Block Design Simplification for our approach.
@@
Retinex-Net [5] 16.77 0.425 0.474 18.37 0.723 17.14 0.756 0.555 587.470
KinD [7] 17.65 0.771 0.175 14.74 0.641 17.28 0.758 8.160 574.950
EnGAN [6] 17.54 0.664 0.326 18.23 0.617 16.49 0.771 114.350 223.430
@@
CPGA-Net+ (BDSF) 22.53 0.812 0.205 20.90 0.800 23.07 0.907 0.020 2.141
TABLE II: The image quality comparison on unpaired data [11, 27, 28, 29, 30] in terms of the NIQE metric, where lower values generally indicate better performance. We represent the first and second ranks with bold and underlined, respectively. For the learning methods, T indicates the traditional approach, U indicates unsupervised learning, and S indicates supervised learning. BDSF means Block Design Simplification for our approach.
Datasets
Original Image and Method Types MEF LIME NPE VV DICM Avg
Low-light Image N/A 4.2650 4.4380 4.3190 3.5350 4.2550 4.1624
NPE [28] T 3.5240 3.9048 3.9530 2.5240 3.7600 3.5332
LIME [11] T 3.7200 4.1550 4.2680 2.4890 3.8460 3.6956
EnlightenGAN [7] U 3.2320 3.7190 4.1130 2.5810 3.5700 3.4430
Zero-DCE [13] U 4.0410 3.7890 3.5041 2.7526 3.1018 3.4377
RUAS [14] U 4.1403 4.2900 4.8713 3.5086 4.5417 4.2704
PairLIE [16] U 4.0862 4.3113 4.0890 3.1595 3.2422 3.7776
CPGA-Net+ [8] U 3.5950 3.2575 3.4438 2.8820 3.0350 3.2427
KinD [7] S 3.8830 3.3430 3.7240 2.3208 2.9888 3.2519
IAT [15] S 3.6188 4.1722 3.2890 2.5270 3.0325 3.3279
FLIGHT-Net [17] S 3.5491 3.7049 3.3311 2.9435 2.8979 3.2853
DDNet [32] S 3.2734 3.4329 3.1135 2.0223 2.6409 2.8970
CPGA-Net [8] S 3.8698 3.7068 3.5476 2.2641 2.6934 3.2163
CPGA-DIA [10] S 3.5880 3.5570 3.1650 2.0930 2.6300 3.0006
CPGA-Net+ S 3.4968 3.0626 3.0886 1.9133 2.8282 2.8779
CPGA-Net+ (BDSF) S 3.4969 3.0655 3.0881 1.9136 2.8268 2.8782
Figure 7: Visual comparison on paired datasets [5].
Figure 8: Visual comparison on unpaired datasets [11, 27, 30].
Figure 9: Visual comparison of unsupervised approaches on unpaired data [28, 29].
IV-D High-Level Vision Task
In this section, we address the challenge of objection detection in low-light environments by utilizing a joint training approach of YOLOv9s [33] with the SOTA approaches of LLIE on the ExDark dataset [31], as illustrated in Table III. Our approach improves the mean Average Precision (mAP) by 0.075 compared to baseline. All the LLIE methods listed here can improve object detection performance, among which the Zero-DCE [13] and our proposed method are the best. The results show that our proposed method can improve not just human perception but machine perception as well.
TABLE III: Comparison of performance metrics between YOLOv9s with CPGA-Net+ and other SOTA methods on the ExDark dataset [31].
@@ IV-D High-Level Vision Task
V Ablation Study
In this section, we analyze the effectiveness of each systematic module and training technique, including the systematic design, the number of Channel-Prior blocks, and loss functions.
V-A Systematic Design and Integration
As shown in Table IV, our method effectively fuses the gamma correction from the global branch to the local branch, resulting in improved overall performance and demonstrating the strength of our approach. By grounding the attention mechanism in gamma correction, we ensure that the enhancement process remains aligned with the non-linearities inherent in both the imaging process and human perception.
On the other hand, we also figured out that the global branch alone performs surprisingly well. In Section III-D, we elaborate on the system underlying the equations and highlight the relationship between $R$ and $R^{\text{GT}}$, and the results suggest that this correlation is stronger than expected. For instance, Table IV(c) and Table IV(e) show only a minor difference in PSNR (0.5 dB), and there is no performance difference between Table IV(e) and Table IV(f) on LOLv1. However, despite the significant contribution of the global branch, Table IV(c) still shows a noticeable gap in achieving optimal performance without incorporating local processing. This underscores the necessity of the local branch to bridge the gap between local and global processing, ensuring the system achieves its highest potential in quality and consistency. Since our pruning approach is based on the concept of block pruning but differs from existing methods [34, 35, 36, 37] and retains its simplicity even without requiring additional training, we named the proposed technique ”Block Design Simplification (BDSF).”
TABLE IV: Ablation study of systematic design. L-G denotes our design of plug-in attention from global to local processing, utilizing the CPGA block to bridge the gap between local and global branches. (f) shares the weights from (e) but performs inference via global processing only, sharing the same design as (c). (e)* denotes using the weights obtained from (e).
@@ V-A Systematic Design and Integration
V-B The Numbers of Channel-Prior Blocks
We explore how varying the number of CP blocks affects the model’s capacity to enhance image quality. The results, summarized in Table V, show that increasing the number of CP blocks leads to an improvement from 0 to 2 blocks but show no significant changes from 2 to 4 blocks. However, both the number of parameters and computational cost (FLOPs) increase with more CP blocks, introducing greater computational demands. Therefore, the optimal number of CP blocks should balance performance gains with resource efficiency. For our final approach, we selected 2 CP blocks to achieve a lightweight and efficient design.
TABLE V: Ablation study of the number of CP blocks. $N=2$ is the default setting of our approach.
@@ V-B The Numbers of Channel-Prior Blocks
V-C Loss Functions
This section examines the impact of various loss function combinations on the model’s performance. We tested L1 loss, Perceptual loss, HDR L1 loss, and SSIM loss, with the results summarized in Table VI. Using the default settings as in CPGA-Net, the combination of L1 and Perceptual losses performs well, yielding a PSNR improvement of 2.82 dB and an SSIM increase of 0.024. The HDR L1 loss significantly enhances all three metrics, with PSNR increasing by 2.86 dB, SSIM by 0.041, and LPIPS decreasing by 0.085. While SSIM loss improves its specific metric with an SSIM boost of 0.031, it is less effective in enhancing PSNR. Ultimately, combining all these losses results in the best overall performance for supervision, which improves PSNR by 4.04 dB, SSIM by 0.083, and LPIPS by 0.118.
TABLE VI: Ablation study of loss functions.
@@ V-C Loss Functions
VI Limitation
While our proposed method shows notable improvements in efficiency and performance, a key limitation warrants consideration. As discussed in the methodology section, our approach benefits from the guidance of Channel-Priors and Gamma Correction, which enhances contrast and visual perception. However, gamma correction for brightening with a small value of estimated gamma can lead to defects or distortions in extreme low-light scenarios, even while maintaining realistic exposure. Fig. 10 shows an example of such a limitation. Addressing this issue will require further refinement for brightening and denoising, which we aim to pursue in future work.
Figure 10: Visualization of extreme low-light scenarios [11, 30]. (a) Original images; (b) Supervised CPGA-Net+; (c) Unsupervised CPGA-Net+. While the enhancement tone appears satisfactory, both methods exhibit defects such as pepper-and-salt noise, resulting in a grainy texture.
VII Conclusion
This work looks deeper at CPGA-Net, utilizing it as an attention mechanism grounded in theoretical formulas. We propose a stacked and modularized attention module to focus on image details. Additionally, we integrate gamma correction into the local branch, creating a Plug-in Attention module for each CP block. This enhancement makes our approach lightweight yet SOTA in performance, maintaining strong efficiency and stable operation across devices with limited computational resources. In the future, we are striving to improve the unsupervised learning of CPGA-Net+ and integrate our approach into HDR imaging and exposure fusion, improving detail preservation in both bright and dark areas by leveraging brightness sensitivity through prior knowledge, such as channel and gamma-correction priors. This will enhance the output’s dynamic range and overall fidelity in real-world applications.
@@ Acknowledgment
This work was supported in part by the Ministry of Science and Technology, Taiwan, under Grant NSTC 113-2221-E-033-055.